OpenStudy (anonymous):

Multiply and simplify: (6x - 11)(6x + 11)

OpenStudy (anonymous):

u know the formula a^2 -b^=(a+b)(a-b)

OpenStudy (jiteshmeghwal9):

use formula :- \(a^2-b^2=(a+b)(a-b)\) Here,a=6x; b=11

OpenStudy (anonymous):

=(6x)^2-(11)^2 =36x^2-121

OpenStudy (anonymous):

ok but then it says to write my answer in descending order,oh sheesh pLease heLp

OpenStudy (anonymous):

on simplification we get (6x-11)(6x+11)